OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. A) y = 4x - 11 B)y = 4x + 11 C)y = 4x + 7 D) y = 4x - 7

OpenStudy (amistre64):

y=mx+b is the general setup for this; m is the slope; they all have a slope of 4

OpenStudy (amistre64):

soo, the next simplest thing to do might be to just use the point values and see whcih one fits

OpenStudy (amistre64):

(-3,-1) is the point; so, x=-3 y=-1 lets try that in one of the options y = 4x + 7 -1 -3 ----------- -1=-12+7 is false so that aint it

OpenStudy (phi):

for each choice, replace x with -3 and y with -1 and do the arithmetic. if it is the right answer, then the right side will equal the left side.

OpenStudy (phi):

amistre did the 3rd choice as an example.